A sample of air has a temperature of 40.0◦C and occupies a volume of 2.30ℓ. At what temperature would it occupy a volume of 6.50ℓ?

1
Expert's answer
2021-12-27T08:01:29-0500

Given:

"t_i=\rm40.0^\circ C"

"V_i=\rm 2.30\: L"

"V_f=\rm 6.50\: L"

"t_f-?"


The constant pressure process equation says

"\frac{T_f}{T_i}=\frac{V_f}{V_i}"

Hence

"T_f=T_i\frac{V_f}{V_i}=(40.0+273)\frac{6.50}{2.30}=885\:\rm K"

"t_f=T_f-273=885-273=612\rm ^\circ C"